# Rethinking Graph Structure Learning in the Era of LLMs

### v1

@@ 4.3 Tree-prompted LLM Inference
Motivation: Based on the structural encoding tree $\mathcal{T}$, we aim to further integrate node language descriptions to enable comprehensive LLM in-context learning, facilitating a thorough understanding of the topology and text insights within the original graph. This approach allows LLMs to perform reliable inference without fine-tuning. By leveraging these inferences, the encoding tree is further optimized to enhance homophily within low-level communities, providing a solid foundation for the subsequent GSL.
Reception-aware Leaf Augmentation. To construct high-quality in-context for LLMs, the quality of the node-specific text is crucial. Therefore, we perform tree-prompted text propagation and aggregation within each low-level community $\mathcal{C}_{i}$ to achieve leaf augmentation for $\epsilon$-guaranteed $\mathrm{T}^{\star}$. The key intuition is that nodes in the low-level communities often exhibit high connectivity and homophily, which facilitates text propagation and aggregation. Based on this, we can filter and combine relevant texts by similarity threshold $\epsilon$ for each leaf node, further mitigating potential noise. The above process is formally defined as:
$\mathrm{t}_{\alpha}^{\star}=\operatorname{Concat}(\mathrm{t}_{\alpha},\{\mathrm{t}_{\beta}|\mathrm{t}_{\beta}\in\mathbb{I}_{\alpha}\}),$ (2)
$\mathrm{sim}(x_{\alpha},x_{\beta})=\frac{\sum_{i=1}^{f\text{-dim}}x_{\alpha}^{i}\cdot x_{\beta}^{i}}{\sqrt{\sum_{i=1}^{f\text{-dim}}(x_{\alpha}^{i})^{2}\sum_{i=1}^{f\text{-dim}}(x_{\beta}^{i})^{2}}},$ (3)
$\mathbb{I}_{\alpha}\!\!=\!\{x_{\beta}\!\mid\!\operatorname{sim}(x_{\alpha},x_{\beta})\!\geq\!\epsilon\},\forall\alpha,\beta\in\operatorname{Set}(\mathcal{C}_{\alpha}),\alpha\!\neq\!\!\beta,$ (4)
@@
$\mathbf{Y}^{cls}=\mathrm{Softmax}(\operatorname{Extract}(\text{Answer})),$ (6)
@@
Leaf Dependency Allocation. Although the structural encoding tree captures the topology insights of the original graph, another critical factor is the text-driven node attribute features. Therefore, we further optimize the low-level communities through leaf dependency allocation, guided by the soft labels derived from the LLM, thereby empowering GSL from the perspective of text insights.
Specifically, we first initialize the leaf-community clusters using the original structural encoding tree. Based on this, we perform adaptive clustering by silhouette coefficient. The core intuition is to maximize homophily within low-level communities, ensuring that nodes within the same community predominantly belong to the same label class. In other words, (1) it considers the majority label classes within each low-level community, reallocating minority-class leaf nodes; and (2) it leverages the silhouette coefficient to adaptively increase clusters, thereby creating new low-level communities that enhance the tree’s ability to reveal potential homophily patterns. A visual example is shown in Fig. 2. At this stage, we have optimized the original structural encoding tree to obtain $\mathcal{T}^{\star}$, which serves as the foundation for the subsequent definition of the tree sampler. Due to space limitations, a detailed description of the above adaptive clustering algorithm can be found in Appendix D.2.
@@ 4.4 Leaf-oriented Two-step Sampling
Motivation. Although structural optimization can be applied to all leaf nodes, we adopt a two-step sampling technology to balance running efficiency and practical performance. Specifically, we first identify a limited set of nodes requiring optimization from a topological perspective. Subsequently, we select a candidate set of nodes closely related to these nodes from a text semantic perspective. Finally, through edge addition or removal, we facilitate training-free GSL. The detailed algorithm can ben found in Appendix D.3
Topology Uncertainty Sampling. As we all know, the core of GSL is to eliminate graph structural noise to enhance data utility. Notably, this noise typically does not pervade the entire graph but exists in certain graph substructures. Therefore, to improve running efficiency, we leverage structural entropy to quantify the priority of nodes in need of structural optimization. Specifically, within each leaf community $\mathcal{C}_{i}$, we use structural entropy $\mathcal{H}^{\mathcal{T}}(\mathcal{G},\alpha)$ to assign probabilities to each leaf node for node-wise sampling:
$\mathcal{P}_{topo}(\alpha)=\frac{\mathrm{exp}(\mathcal{H}^{\mathcal{\mathcal{T}}}(\mathcal{G},\alpha))}{{\underset{\gamma\in\mathcal{C}_{\alpha}}{\sum}}{\mathrm{exp}(H^{\mathcal{\mathcal{T}}}(\mathcal{G},\gamma))}}.$ (7)
@@
$\mathcal{P}_{sema}^{\alpha}(\beta)=\frac{\exp(\operatorname{sim}(y^{cls}_{\beta},{y}^{cls}_{\alpha}))}{{\underset{\phi\in\mathcal{C}_{\mathcal{H}},\phi\neq\alpha}{\sum}}{\mathrm{exp}(\mathrm{sim}(y^{cls}_{\phi},y^{cls}_{\alpha}))}}.$ (8)
@@ 5 Experiment
Table 1: Node classification accuracy(%) on 10 TAG datasets. Shown is the mean of 10 runs with different random seeds. Highlighted are the top first, second, and third accuracy. OOM denotes out of memory and OOT denotes time limits of 72 hours exceeded.
@@ 5.1 Experiment Setup
Datasets. We evaluate LLaTA and other baselines on 10 widely adopted TAG datasets across multiple domains, including 3 citation networks (Chen et al., 2024), 1 knowledge network (Chen et al., 2024), 2 social networks (Huang et al., 2024), and 4 e-commerce networks (Yan et al., 2023). Among these, the Ratings and Child exhibit heterophily, while the remaining datasets follow homophily. Further details on these datasets can be found in Appendix E.
Baselines. We compare LLaTA with 5 coupled GSL methods (Chen et al., 2020; Fatemi et al., 2021; Zhao et al., 2021; Wu et al., 2023; Zou et al., 2023), 6 decoupled GSL methods (Jin et al., 2020; Liu et al., 2022b; Li et al., 2022; Liu et al., 2022a; Nguyen et al., 2023; Bi et al., 2024), and 2 LLM-based GSL methods(Guo et al., 2024; Zhang et al., 2024). Based on this, we apply prevalent GCN (Kipf & Welling, 2016), GAT (Veličković et al., 2017), and GraphSAGE (Hamilton et al., 2017) as the downstream backbone. Details of the above baselines can be found in Appendix F.
[…]
Table 2: Comparison of node clustering accuracy(%).
@@ 5.2 Performance Comparison
Node Clustering. To further answer Q1, we conducted a comprehensive comparison of LLaTA with GSL methods that have demonstrated strong performance in node classification, extending to node clustering. As shown in Table 2, LLaTA consistently outperforms competing methods across all four datasets, significantly enhancing the effectiveness of GNNs in unsupervised tasks. These results confirm that our approach maintains strong performance across diverse tasks, demonstrating its generalization and adaptability.
@@ 5.3 Ablation Study
To answer Q2, we conduct an ablation study shown in Table 3 and provide a case study in Appendix G, which offers a detailed analysis of our method’s complete pipeline by tracing and visualizing the data flow. Meanwhile, we provide LLM backbone analysis in Appendix H. In the ablation study, we use TO to denote tree optimization, while TOLLM and SamLLM denote the tree optimization and sampling processes. For the ablation setup, the LLM inference results are replaced with the initial features. Additionally, LLMNE and LLMRW represent LLM inference with topology-aware tree in-context information derived from 1-hop neighbors and random walk sequences, rather than the tree, respectively.
From the ablation study, we obtain the following key conclusions: (1) Removing TO leads to a significant performance decline, indicating that the tree plays a crucial role in LLM-based GSL. (2) Replacing LLM inference results with initial node features to guide TO leads to a performance decline. This highlights the critical role of LLM-generated contextual information, as initial node features alone are inadequate for capturing complex semantic relationships. (3) Utilizing 1-hop neighbors and random walk sequences to provide topology-aware in-context information also results in performance degradation. This highlights the necessity of the SE-based hierarchical structural encoding tree. Meanwhile, SE provides valuable guidance for node sampling, a capability that cannot be effectively achieved solely with 1-hop neighbors or random walk sequences.
Table 3: Ablation study of LLaTA on three TAG datasets.
Table 4: Improvements of LLaTA with different GNN backbones.
@@
5.4 Robustness Analysis
To answer Q3, we conduct a thorough analysis of the LLaTA’s robustness from the following three aspects:
Downstream Backbones. According to Table 4, LLaTA consistently outperforms the base models across all datasets, achieving notable improvements of up to 10.18% for GCN and 9.31% for GAT. Although the performance gains for GraphSAGE are relatively smaller, this can be attributed to its strong inherent node representation capability. In a nutshell, these results demonstrate LLaTA’s robustness in enhancing various GNN backbones, highlighting its broad applicability across mainstream GNN architectures.
Figure 3: Node classification performance of LLaTA with different $K$, $\theta$ and $r$. $\theta$ and $r$ are analyzed on the Child dataset.
Figure 4: Node classification performance on the Cora dataset under real-world scenarios (sparse [left] and noise [right]).
Hyperparameter. We conducted a comprehensive analysis of the four key hyperparameters of LLaTA: $K,\epsilon,\theta$, and $r$. The experimental results for $K,\theta$, and $r$ are presented in Fig. 3, while the analysis of $\epsilon$ is provided in Appendix I. In Fig. 3 (left), we observe that the optimal encoding tree height $K$ varies across different datasets. Specifically, the optimal value is $K=3$ for WikiCS and Pubmed, and $K=2$ for Instagram. However, when $K$ increases to 5, the performance of LLaTA significantly declines across all three datasets, indicating that excessive tree depth may lead to over-fitting or loss of structural information. In Fig. 3 (right), we analyze LLaTA’s sensitivity to the candidate set size $\theta$ in semantic similarity sampling and the two-step sampling frequency $r$ on the Child dataset. The results indicate that an $r$ value between 5 and 25 yields higher accuracy, corresponding to an appropriately improved graph density. In contrast, LLaTA is less sensitive to the choice of $\theta$, achieving stable performance within the range of $\theta\in[5,15]$. A smaller $\theta$ may result in insufficient semantic similarity sampling, whereas a larger $\theta$ could introduce noise into the candidate set, negatively impacting performance.
Figure 5: Comparison of training and inference time.
@@ 5.5 Efficiency analysis
Considering the time complexity of SEGSL (Zou et al., 2023) $O(n^{2}+nlog^{2}n+n)$, we analyze the theoretical algorithm complexity of LLaTA. The overall time complexity of LLaTA is $O(nlog^{2}n+nm^{2}+nt_{LLM}+n)$, where $n$ represents the number of nodes, and $m$ denotes the size of low-level communities. Specifically, the time complexity for encoding tree construction, as discussed in Sec. 4.2, is $O(nlog^{2}n)$. The complexity of tree-prompted LLM inference, as described in Sec. 4.3, is $O(nm^{2}+nt_{LLM})$, where $t_{LLM}$ refers to the time required for a single LLM inference. Additionally, the time complexity of the two-step sampling process, outlined in Sec. 4.4, is $O(\theta n)$.
@@ 6 Conclusion
In this paper, we first introduce a novel optimization framework for GSL by rethinking its integration with LLMs, addressing the challenges of incorporating textual information. Subsequently, We propose LLaTA as an instantiation of this novel framework, which leverages a structural encoding tree to achieve efficient LLM in-context learning. This enables the LLM to comprehensively understand both topology and text insights from the original graph, ultimately relying on reliable inference to obtain improved graph structure. To this end, LLaTA eliminates the need for computation-intensive fine-tuning and achieves SOTA performance. Inspired by the noise experiments, a promising direction is to design more robust tree construction algorithms, providing a solid foundation for subsequent LLM-empowered GSL.

### v2

@@ 4.3 Tree-prompted LLM Inference
Motivation: Based on the structural encoding tree $\mathcal{T}$, we further integrate node language descriptions to enable comprehensive in-context learning with LLMs, fostering a better understanding of both topology and textual information in the original graph. This allows LLMs to make reliable inferences without fine-tuning. These inferences are then used to optimize the encoding tree, enhancing homophily within low-level communities and laying a strong foundation for subsequent GSL.
Reception-aware Leaf Augmentation. To construct effective in-context prompts for LLMs, the quality of node-specific text is essential. Thus, we perform tree-guided text propagation and aggregation within each low-level community $\mathcal{C}^{\ell}(i)$ to achieve leaf augmentation for $\epsilon$-guaranteed $t^{\star}$. The key idea is that nodes in low-level communities typically show high connectivity and homophily, enabling efficient text sharing. Using this property, we filter and combine relevant texts for each leaf node based on a similarity threshold $\epsilon$, helping to reduce noise. This process is formally defined as:
$t^{\star}_{\alpha}=\operatorname{Concat}\left(\left\{t_{\alpha}\}\cup\{t_{j}\mid j\in\text{Top}_{k}\left(\left\{w_{\alpha\beta}\mid\beta\in\mathcal{C}^{\ell}(\alpha)\right\}\right),w_{\alpha\beta}\geq\epsilon\right\}\right),$ (2)
$w_{\alpha\beta}=\frac{\exp\left(\operatorname{sim}(x_{\alpha},x_{\beta})\right)}{\sum\limits_{\gamma\in\mathcal{C}^{l}(\alpha)}\exp\left(\operatorname{sim}(x_{\alpha},x_{\gamma})\right)},\>\>\>\>\operatorname{sim}(x_{\alpha},x_{\beta})=\frac{\sum_{i=1}^{f\text{-dim}}x_{\alpha}^{i}\cdot x_{\beta}^{i}}{\sqrt{\sum_{i=1}^{f\text{-dim}}(x_{\alpha}^{i})^{2}\sum_{i=1}^{f\text{-dim}}(x_{\beta}^{i})^{2}}}.$ (3)
@@
$z_{\alpha}=\operatorname{Extract}(ANS_{\alpha}),\>\>\>\>\text{ANS}_{\alpha}=\text{LLM-}\operatorname{CoT}(\mathcal{D}_{task},\>\mathcal{D}_{topo},\>\mathcal{D}_{sema},\>t^{\star}_{\alpha}),$ (4)
@@
Theorem 3 (Error Bound Between Soft labels and True Labels).
Given two leaf nodes $\alpha$ and $\beta$ in $\mathcal{T}$, the error between soft label similarity and true label similarity is bounded by:
$\varepsilon^{y}(\alpha\beta)=|\operatorname{sim}(y^{cls}_{\alpha},y^{cls}_{\beta})-\operatorname{sim}(y_{\alpha},y_{\beta})|\leq\delta\cdot(1-\epsilon),$ where $\delta$ is a constant that depends on the LLM’s in-context learning ability and $y_{\alpha}$ is the true label of $\alpha$. (The proof is detailed in Appendix F.3.)
Based on Theorem 3, we have that the similarity of soft labels $y^{cls}$ can approximate the similarity of ground-truth labels $y$ with controllable bias $\varepsilon^{y}(\alpha\beta)$. This provides robust theoretical guarantee and support for the subsequent Leaf Dependency Allocation and Semantic Similarity Sampling.
Leaf Dependency Allocation. Although the structural encoding tree captures the topology of the original graph, another critical factor is the text-driven node attributes. Therefore, we optimize the low-level communities by reallocating leaf dependencies, guided by soft labels from the LLM, thus enhancing GSL with text insights. To achieve this, we first initialize the leaf-community clusters based on the original structural encoding tree. Next, we apply adaptive clustering using the silhouette coefficient to refine community assignments. The main goal is to enhance homophily within low-level communities, ensuring that most nodes in each cluster share the same label. Specifically, (1) we reallocate leaf nodes with minority labels to align with the dominant class in each community, and (2) we adaptively increase the number of clusters by leveraging the silhouette coefficient, allowing the formation of new communities that better capture homophily. An illustrative example is shown in Fig.2. This process yields the optimized structural encoding tree $\mathcal{T}^{\star}$, which serves as the foundation for defining the tree sampler. The adaptive clustering algorithm is detailed in Appendix E.2.
@@ 4.4 Leaf-oriented Two-step Sampling
Motivation. Although structural optimization can be applied to all leaf nodes, we adopt a two-step sampling technology to balance running efficiency and practical performance. Specifically, we first identify a limited set of nodes requiring optimization from a topological perspective. Subsequently, we select a candidate set of nodes closely related to these nodes from a text semantic perspective. Finally, through edge addition or removal, we facilitate training-free GSL. The detailed algorithm of the two-step sampling can ben found in Appendix E.3.
Theorem 4 (High-Entropy Nodes Require Supervision).
In a structural encoding tree $\mathcal{T}$ constructed via entropy minimization, nodes with higher structural entropy $\mathcal{H}^{\mathcal{T}}(\mathcal{G},\alpha)$ indicate: higher topological uncertainty within their local structural context. (The proof is detailed in Appendix F.4.)
Topology Uncertainty Sampling. As we all know, the core of GSL is to eliminate graph structural noise to enhance data utility. Notably, this noise typically does not pervade the entire graph but exists in certain graph substructures. According to Theorem 4, nodes with higher structural entropy are more likely to reside in topologically uncertain regions and thus benefit more from supervision-aware structural refinement. Therefore, to improve running efficiency, we leverage structural entropy to quantify the priority of nodes in need of structural optimization. Specifically, within each leaf community $\mathcal{C}^{\ell}(i)$, we use structural entropy $\mathcal{H}^{\mathcal{T}}(\mathcal{G},\alpha)$ to assign probabilities to each leaf node for node-wise sampling:
$\mathcal{P}_{topo}(\alpha)=\frac{\mathrm{exp}(\mathcal{H}^{\mathcal{\mathcal{T}}}(\mathcal{G},\alpha))}{{\underset{\gamma\in\mathcal{C}^{\ell}(\alpha)}{\sum}}{\mathrm{exp}(H^{\mathcal{\mathcal{T}}}(\mathcal{G},\gamma))}}.$ (6)
@@
$\mathcal{P}_{sema}^{\alpha}(\beta)=\frac{\exp(\operatorname{sim}(y^{cls}_{\beta},{y}^{cls}_{\alpha}))}{{\underset{\gamma\in\mathcal{C}^{\ell}(\alpha),\gamma\neq\alpha}{\sum}}\!\!\!{\mathrm{exp}(\mathrm{sim}(y^{cls}_{\gamma},y^{cls}_{\alpha}))}}.$ (7)
@@ 5.1 Experiment Setup
We evaluate LLaTA and 13 GSL baselines on 10 widely adopted TAG datasets across multiple domains. The details on these datasets and baselines can be found in Appendix G-H. Due to space limitations, the hyperparameter settings used in the experiments are provided in the Appendix I.
@@ 5.2 Performance Comparison
Node Clustering. To further answer Q1, we conducted a comprehensive comparison of LLaTA with GSL methods that have demonstrated strong performance in node classification, extending to node clustering. As shown in Table 3, LLaTA consistently outperforms competing methods across all four datasets, significantly enhancing the effectiveness of GNNs in unsupervised tasks. These results confirm that LLaTA is applicable to diverse tasks, demonstrating its generalization and adaptability.
Table 1: Node classification accuracy(%) on 10 TAG datasets. Shown is the $mean\pm std$ of 10 runs with different random seeds. Highlighted are the top first, second, and third accuracy. OOM denotes out of memory and OOT denotes time limits of 72 hours exceeded.
Table 2: Comparison of node clustering.
Table 3: Ablation study result of LLaTA.
@@ 5.3 Ablation Study
To answer Q2, we conduct an ablation study shown in Table 3 and provide a case study in Appendix J, which provides a detailed analysis of our method’s pipeline by visualizing the data flow. Meanwhile, we provide LLM backbone analysis in Appendix K. In the ablation study, we use TO to denote tree optimization, while TOLLM and SamLLM denote the tree optimization and sampling processes with LLM. For the ablation setup, the LLM inference results are replaced by the initial features. Additionally, LLMNE and LLMRW represent LLM inference with topology-aware tree in-context information derived from 1-hop neighbors and random walk sequences, rather than the tree, respectively.
From the ablation study, we obtain the following key conclusions: (1) Removing TO leads to a significant performance decline, indicating that the tree plays a crucial role in LLM-based GSL. (2) Replacing LLM inference results with initial node features to guide TO and Sam leads to a performance decline. This highlights the critical role of LLM-generated contextual information, as initial node features alone are inadequate for capturing complex semantic relationships. (3) Utilizing 1-hop neighbors and random walk sequences to provide topology-aware in-context information also results in performance degradation. This highlights the necessity of the SE-based hierarchical structural encoding tree. Meanwhile, SE provides valuable guidance for node sampling, a capability that cannot be effectively achieved solely with 1-hop neighbors or random walk sequences.
5.4 Robustness Analysis
To answer Q3, we conduct a thorough analysis of the LLaTA’s robustness from the following aspects:
Downstream Backbones. According to Table 5, LLaTA consistently outperforms the base models across all datasets, achieving notable improvements of up to 10.18% for GCN and 9.31% for GAT. Although the gains for GraphSAGE are smaller, this can be attributed to its strong inherent node representation capability. Overall, these results demonstrate LLaTA’s robustness in enhancing various GNN backbones, highlighting its broad applicability across mainstream architectures.
Table 4: Improvements (compare to GNN) of LLaTA with different GNN backbones.
@@ 5.4 Robustness Analysis
Table 5: Time complexity analysis of LLaTA and other GSL method.
Method Time complexity SEGSL $\mathcal{O}(n^{2}+nlog^{2}n+n)$ GraphEdit $\mathcal{O}(pt_{\text{LLM-r}}+qd^{2}+n^{2}+(m+nk)t_{\text{LLM-i}})$ LLM4RGNN $\mathcal{O}(pt_{\text{LLM-r}}+mt_{\text{LLM-i}}+qd^{2}+n^{2})$ LLaTA $O(nlog^{2}n+n|\mathcal{C}|^{2}+nt_{\text{LLM-i}}+n)$
Figure 3: Node classification performance on the Cora and Citeseer dataset under real-world scenarios (sparse [Edge Removal] and noise [Edge Addition]).
@@
Hyperparameter. We conducted a comprehensive analysis of the four key hyperparameters of LLaTA: $K,\epsilon,\theta$, and $r$. The experimental results for $K,\theta$, and $r$ are presented in Fig. 5, while the analysis of $\epsilon$ and the experimental results of $\theta,r$ on other datasets are provided in Appendix L. In Fig. 5 (left), we see that the optimal encoding tree height $K$ varies by dataset, with $K=3$ for WikiCS and Pubmed, and $K=2$ for Instagram. However, when $K$ increases to 5, performance significantly drops across all datasets, suggesting that excessive tree depth may cause overfitting or loss of structural information. In Fig. 5 (right), we analyze LLaTA’s sensitivity to the candidate set size $\theta$ in semantic similarity sampling and the two-step sampling frequency $r$ on the Cora dataset. The results indicate that an $r$ value between 3 and 10 yields higher accuracy, corresponding to an appropriately improved graph density. In contrast, LLaTA is less sensitive to the choice of $\theta$, achieving stable performance within the range of $\theta\in[5,15]$. A smaller $\theta$ may result in insufficient semantic similarity sampling, whereas a larger $\theta$ could introduce noise into the candidate set, negatively impacting performance.
Figure 4: Hyperparameter analysis of $K$, $\theta$ and $r$. $\theta$ and $r$ are analyzed on the Cora dataset.
Figure 5: Comparison of training and inference time for LLM-based GSL method.
@@ 5.5 Efficiency analysis
In Table 5, we analyze the theoretical time complexity of LLaTA and existing LLM-based GSL methods. Where $q$ and $p$ is the number of samples for fine-tuning and edge predictor training, respectively. $t_{\text{LLM-r}}$ and $t_{\text{LLM-i}}$ denote the fine-tuning and inference time of LLM, respectively. The magnitude of $t_{\text{LLM-i}}$ should be adaptively determined according to the specific requirements of different inference tasks. For LLaTA, The time complexity can be divided into three parts: (a) The complexity for tree construction in Sec. 4.2 is $\mathcal{O}(nlog^{2}n)$, (b) The complexity of tree-prompted LLM inference in Sec. 4.3 is $\mathcal{O}(n|\mathcal{C}|^{2}+nt_{\text{LLM-i}})$, (c) The complexity of the two-step sampling in Sec. 4.4 is $\mathcal{O}(\theta n)$. Complete runtime results of LLaTA across all datasets are provided Appendix M.
@@ 6 Conclusion
In this paper, we first introduce a novel optimization framework for GSL by rethinking its integration with LLMs, addressing the challenges of incorporating textual information. Subsequently, We propose LLaTA as an instantiation of this novel framework, which leverages a structural encoding tree to achieve efficient LLM in-context learning. This enables the LLM to comprehensively understand both topology and text insights from the original graph, ultimately relying on reliable inference to obtain improved graph structure. To this end, LLaTA eliminates the need for costly fine-tuning and achieves SOTA performance. Inspired by the noise experiments, a promising direction is developing more robust tree optimization algorithms, laying a solid foundation for LLM-empowered GSL.